OpenStudy (goformit100):

The denominator of a rational number is greater than its numerator by 8. If the numerator is increased by 17 and the denominator is decreased by 1, the number obtained is 3/2. Find the rational number.

OpenStudy (goformit100):

@Justin9529

OpenStudy (justin9529):

if the numerator=x then the denominator=x+8 the fraction=x/x+8 if numerator is increased by 17&denominator is decreased by 1then the fraction is x+17/ x+8-1=x+17/ x+7 x+17/ x+8=3/2 cross multiplication =2(x+17)=3(x+8) =2x+34=3x+24 3x-2x=34-24 x=10 numerator=x=10 denominator=x+8=18